OpenStudy (anonymous):

Solve the following system of equations. x + 3y – z = 2 x – 2y + 3z = 7 x + 2y – 5z = –21 (2, 3, 5) (–2, 3, 5) (2, –3, 5) (2, 3, –5)

OpenStudy (anonymous):

it can be solved either by substitution or matrix method

OpenStudy (anonymous):

pretty lengthy process

OpenStudy (texaschic101):

x + 3y - z = 2 x - 2y + 3z = 7 x + 2y - 5z = - 21 ----------------- start with the first two equations.... x + 3y - z = 2 --->(-1)x + 3y - z = 2 x - 2y + 3z = 7 --------------- -x - 3y + z = -2 (result of multiplying by -1) x - 2y + 3z = 7 ---------------add -5y + 4z = 5 now take the last two equations and eliminate x x - 2y + 3z = 7 x + 2y - 5z = - 21 -->(-1)x + 2y - 5z = - 21 ----------------- x - 2y + 3z = 7 -x - 2y + 5z = 21 (result of multiplying by -1) ------------------add - 4y + 8z = 28 now add the answers you got from the first two equations and the last two equations... -5y + 4z = 5 --> (-2)-5y + 4z = 5 -4y + 8z = 28 ------------- 10y - 8z = - 10 (result of multiplying by -2) -4y + 8z = 28 --------------add 6y = 18 y = 18/6 y = 3 now sub 3 in for y in either of the answers you got from the first two equations or the last two equations... -4y + 8z = 28 -4(3) + 8z = 28 -12 + 8z = 28 8z = 28 + 12 8z = 40 z = 40/8 z = 5 now sub 3 in for y and 5 in for z in any of the original equations... x + 3y - z = 2 x + 3(3) - 5 = 2 x + 9 - 5 = 2 x = 2 + 5 - 9 x = 7 - 9 x = -2 check... x - 2y + 3z = 7 -2 - 2(3) + 3(5) = 7 -2 - 6 + 15 = 7 - 8 + 15 = 7 7 = 7 (correct) ANSWER : x = -2, y = 3, z = 5 (-2,3,5)